The perimeter of a rectangle is 76cm. The length of the rectangle is 2 cm more than twice he width. Find the dimensions of the rectangle.
Answer by Adrish(44)

Solutions: Let the width be "x"
so the length will be 2+2x
Given: Perimeter = 76cm
P=2L + 2W
76 = 2(2+2x) + 2x
76 = 4 + 4x +2x
76 = 4 + 6x
76-4 = 4-4 + 6x
72 = 6x
x= 72/6
x= 12 cm
Therefore the length= 2 + 2(12)
= 26 cm
and Width = 12 cm